# Radial Basis Function Techniques for Neural Field Models on Surfaces

### v1

@@ Abstract
We present a numerical framework for solving neural field equations on surfaces using Radial Basis Function (RBF) interpolation and quadrature. Neural field models describe the evolution of macroscopic brain activity, but modeling studies often overlook the complex geometry of curved cortical domains. Traditional numerical methods, such as finite element or spectral methods, can be computationally expensive and challenging to implement on irregular domains. In contrast, RBF-based methods provide a flexible alternative by offering interpolation and quadrature schemes that efficiently handle arbitrary geometries with high-order accuracy. We first develop an RBF-based interpolatory projection framework for neural field models on general surfaces. Quadrature for both flat and curved domains are derived in detail, ensuring high-order accuracy and stability as they depend on RBF hyperparameters (basis functions, augmenting polynomials, and stencil size). Through numerical experiments, we demonstrate the convergence of our method, highlighting its advantages over traditional approaches in terms of flexibility and accuracy. We conclude with an exposition of numerical simulations of spatiotemporal activity on complex surfaces, illustrating the method’s ability to capture complex wave propagation patterns.
@@ 1 Introduction
Neural field equations are nonlinear integro-differential equations used as large-scale models of neuronal activity, providing tractable equations amenable to analysis and simulations on a diversity of domains [64, 1, 10]. Large-scale neuronal activity is strongly determined by the architecture of synaptic connections between neurons [17]. Associated models typically grapple with the high volume of neurons and synapses, especially in primate cortex, with approximations such as coarsening, homogenization, and timescale separations [28, 8]. Neural fields describe cortical domains as continuum excitable media and their network connectivity as a weight kernel within a nonlinear integral operator, allowing us to apply theory from nonlinear waves and other methods from partial differential equations (PDE) [10]. This kernel defines a spatially dependent coupling function that assigns synaptic strength based on the locations of pre- and post-synaptic neurons. Most analyses and simulations of neural fields therefore focus on the treatment and effects of this nonlinear and nonlocal interaction term as it conveys local dynamics throughout the domain.
@@
The derivation of neural fields is grounded in the known spatial and functional structure of synaptic connectivity throughout cerebral cortex, justifying the typical spatial coarse graining required to attain the standard model [65, 57]. For the convenience of analysis and simulation, most modeling studies consider spatial domains and connectivity functions that take canonical forms (e.g., distance-dependent weight kernels on planar domains) [18]. This has yielded considerable dividends in the understanding of the wide diversity of spatiotemporal dynamics these models can support, as well as novel theory for the stability of solutions to nonlocal equations. In seminal work, Amari (1977) simplified the nonlinearity appearing on the integral as a Heaviside step function [1], allowing for the implementation of interface methods and direct construction of traveling wave solutions [44, 20]. Cleverly chosen weight kernels with rational Fourier transforms allow for the conversion of nonlocal neural fields to local PDE for more efficient analysis and simulation [37, 21]. However, tractable models often overlook more complex geometries. Here we advance a method that is highly flexible to the formulation of neural fields on arbitrary domains with curvature and nonstandard weight kernels.
Realistic observable cortical domains are often two dimensional surfaces with curvature embedded in three dimensional space. While the cortex itself is of course three-dimensional, imaging studies most commonly access neural activity on the surface via functional magnetic resonance or voltage sensitive dye imaging [38, 31]. Modeling efforts in this direction have focused on prototypical curved domains such as the sphere [52, 12, 62] or the torus [35]. Solutions on such domains can then be conveniently expressed using spectral methods with clean and complete orthogonal eigenbases, but complex domains, that are not expressed as Cartesian products, cannot leverage this approach. Recently, there has been interest in simulating neural fields on arbitrary smooth surfaces [41], but so far the associated algorithms have no theoretical convergence guarantees, and empirical error measurements suggest the method is only accurate to first order.
A framework for assessing convergence for any projection-based numerical neural-field solver was recently developed in [4]. This approach decomposes numerical error in the solution into three main components: the error made by the spatial projection, the error from numerical time-integration, and the error from the quadrature. Importantly, the paper provides a unified operator-theoretic formulation that enables rigorous convergence analysis for Galerkin and collocation schemes. This analysis not only clarifies how the error scales with spatial resolution, but also guides the development of more efficient discretizations for neural field simulations on complex geometries.
In the current work, we present a high-order numerical method for neural field simulations, and partially address theory regarding numerical convergence within the projection method framework. This high-order accuracy is achieved using radial basis function formulae for quadrature (RBF-QF) [49, 51, 50] and radial basis function (RBF) interpolation as the projector. High-order methods are essential to allow for simulations of neural field equations onto complex surfaces, as they allow for high accuracy for fewer computational resources. Specifically, approximating the integral operator using a quadrature on a mesh with $n$ points, requires $O(n^{2})$ operations [^1] For certain domains, one can leverage the FFT to achieve $\mathcal{O}(n\log n)$ complexity, but these methods do not generalize to arbitrary surfaces.. In Section 2 we describe the mathematical framework for collocation and projection schemes for neural field models, formulate the associated discrete ODEs, and state convergence results for projection schemes. In Section 3 we review the RBF-QF algorithm first for flat domains, then for smooth closed manifolds, highlighting the role of local interpolation and the construction of projection-based quadrature on curved geometries. In Section 4 we conduct several numerical experiments verifying convergence rates for RBF-QF and our new neural field solver. In Section 5 we demonstrate the power of this method by simulating several interesting neural field systems on non-trivial domains.
@@ 2 Collocation schemes for the neural field model
Step 1: collocation. Select $n$ points $\Xi=\{\bm{x}_{i}:i\in\mathbb{N}_{n}\}\subset D$ in the domain, where $n$ may depend on a discretization parameter $h$ so $n(h)\to\infty$ as $h\to 0$. We impose that eq. 1 holds at each node, that is, we collocate the equation on $\Xi$. This leads to a functional equation relating $\{\partial_{t}u(t,\bm{x}_{i})\}_{i}$ and $\{u(t,\bm{x}_{i})\}_{i}$ to the function $u(t,\underline{\phantom{n}})$, appearing under the integral of eq. 1.
@@
2.1 Interpolatory projection
In this setup, one seeks a solution $u$ to the neural field equations as a mapping on $[0,T]$ to $\mathbb{X}=C(D)$, the space of continuous functions on $D$. An interpolatory projector $P_{n}$ is a linearly projects any function $v\in\mathbb{X}$ to a function $v_{n}=P_{n}v$ in an $n$-dimensional subspace $\mathbb{X}_{n}=\operatorname{Span}\{L_{1},\ldots,L_{n}\}\subset\mathbb{X}$, with the property
$P_{n}\colon\mathbb{X}\to\mathbb{X}_{n},\qquad P_{n}v=v,\qquad\text{on $\Xi$.}$ (2)
The approximant then takes the form
$v_{n}(\bm{x}):=(P_{n}v)(\bm{x})=\sum_{j=1}^{n}v(\bm{x}_{j})L_{j}(\bm{x}),\qquad x\in D,$
and a useful characterization of interpolatory projectors is the following (See [2, 3] and [4, adapted from Proposition 4.1]):
Proposition 2.1 (Characterization of interpolatory projectors).
If a projection operator $P_{n}\colon\mathbb{X}\to\mathbb{X}_{n}$ satisfies the interpolation condition eq. 2, then for any $r\in\mathbb{X}$
$P_{n}r=0\quad\text{on $D$}\qquad\text{if and only if,}\qquad P_{n}r=0\quad\text{on $\Xi$}.$ (3)
This property is useful when the function $r$ is a residual error, as we will see in a moment.
2.2 General hypotheses
In order to formulate the neural field problem in operator form, we make the following standing assumptions.
Hypothesis 2.1 (General hypotheses).
@@ 2.2 General hypotheses
As shown in [4], Hypothesis 2.1 provides a natural functional setup for the neural field problem posed on $\mathbb{X}=C(D)$, minimally extending [48] to compact cortices $D$ and for studying numerical schemes. It is possible to deal with more generic kernels than the one considered here, but we restrict our work to continuous kernels for simplicity of exposition. Frameworks in [48, 4] could be leveraged for such extensions.
2.3 Neural field in operator form
To set the stage for our collocation scheme, we first rewrite the neural field eq. 1 in operator form
$u^{\prime}(t)=-u(t)+WF(u(t))+g(t)=:N(t,u(t)),\qquad t\in J\equiv[0,T],$ (4)
@@ 2.3 Neural field in operator form
[…]
Note, eq. 1 evolves the $\mathbb{R}$-valued function $u\colon D\times J\to\mathbb{R}$, whereas eq. 4 evolves the $\mathbb{X}$-valued function $U\colon J\to\mathbb{X}$, $t\mapsto u(\underline{\phantom{n}},t)$. With a small abuse of notation, we adopt the same symbol for both functions ($u$), as for the forcing function $g$.
eq. 4 describes a Cauchy problem on the Banach space $\mathbb{X}$, which is well-posed under Hypothesis 2.1. That is, there exists a unique $u\in C^{1}(J,\mathbb{X})$ satisfying eq. 4, as shown in [4, Lemma 2.7] (See [48] for analogous results on unbounded cortices).
2.4 Step 1: collocation scheme
@@
In view of proposition 2.1, this is equivalent to imposing that eq. 1 holds at all $\bm{x}_{i}\in\Xi$ (hence we have collocated the equation, and completed Step 1). That is, we require the eq. 1 to hold exactly at the collocation points $\{\bm{x}_{i}\}_{i=1}^{n}$, reducing the infinite-dimensional problem to a finite-dimensional system of ODEs. In [4, Theorem 3.1] it is shown that, under Hypothesis 2.1, also the projected problem is well-posed, admitting a unique solution $u_{n}\in C^{1}(J,\mathbb{X}_{n})$.
The projected eq. 6 evolves the function $u_{n}(t)$ in $\mathbb{X}_{n}$. This form is useful in the analysis of the scheme, but for implementations we pursue an ODE approximating the $n$ coefficients $\alpha_{i}(t):=u_{n}(t)(\bm{x}_{i})$ using eq. 3, which lets us pass from eq. 6 to an equivalent ODE in $\mathbb{R}^{n}$. We define the residual
$R\colon C^{1}(J,\mathbb{X})\to C(J,\mathbb{X}),\qquad u\mapsto u^{\prime}-N(\underline{\phantom{n}},u),$
and rewrite eq. 6 as
$P_{n}\bigl{(}R(u_{n})(t)\bigr{)}=0,\qquad t\in J,\qquad P_{n}\bigl{(}u_{n}(0)-v\bigr{)}=0,$
which, by property eq. 3 is equivalent to
@@
Since $u_{n}(t)\in\mathbb{X}_{n}$, we set $u_{n}(t)=\sum_{j\in\mathbb{N}_{n}}\alpha_{j}(t)L_{j}$, exploit the Lagrange property of the basis $\{L_{j}\}$, and obtain an ODE in $\mathbb{R}^{n}$
$\alpha^{\prime}(t)=-\alpha(t)+K(\alpha(t))+\gamma(t)\qquad t\in J,$ (7)
@@
[…]
@@ 2.5 Step 2: quadrature and Discrete Projection Schemes
ODE eq. 7 on $\mathbb{R}^{n}$ is not yet implementable on a computer, because generally the integrals in $K$ can not be evaluated in closed form, and must be approximated by numerical quadrature. Such an approach generates a computationally implementable algorithm using Discrete Projection Schemes.
@@
$W_{n}\colon\mathbb{X}\to\mathbb{X}_{n},\qquad(W_{n}v)(\bm{x})=Q_{n}(w(\bm{x},\underline{\phantom{n}})v),$ (8)
@@
$\tilde{u}^{\prime}_{n}(t)=-\tilde{u}_{n}(t)+P_{n}W_{n}F(\tilde{u}_{n}(t))+P_{n}g(t),\qquad t\in J,$ (9)
@@
Following similar steps to Section 2.4, we use proposition 2.1 to obtain
$\tilde{\alpha}^{\prime}(t)=-\tilde{\alpha}(t)+\tilde{K}(\tilde{\alpha}(t))+\gamma(t)\qquad t\in J,$ (10)
@@ 2.6 Convergence of the collocation scheme
Under Hypothesis 2.1, if $P_{n}v\to v$ for all $v\in\mathbb{X}$, then $u_{n}\to u$ in $C(J,\mathbb{X})$. Further, there exist positive constants $m$ and $M$, independent of $n$, such that
@@ 3.1 Local Interpolation
The first use of RBF interpolation was by Hardy in 1971 [30] to construct a topography from scattered elevation measurements. The interpolant was a linear combination of radially-symmetric basis functions called Multiquadrics. Critically, the centers of these basis functions coincided with the locations of the measurements, and as a result, the interpolants were guaranteed to exist and be unique. The Mairhuber-Curtis theorem shows that such guarantees break down when basis functions are chosen independently of the interpolation nodes [23, 40, 22] as occurs in multivariate polynomial interpolation. Later developments showed that other radially symmetric basis functions could be used similarly, with resulting interpolants exhibiting spectral convergence as the number of nodes increased  [14, 67]. For a comprehensive introduction to RBF interpolation, see Fasshauer (2007) [23].
The price of spectral convergence, however, is that finding interpolant coefficients requires solving an $n\times n$ linear system, which becomes prohibitively expensive and ill-conditioned as $n$ (the number of interpolation nodes) increases. There are two approaches to remedy this limitation. First, one can choose compactly supported basis functions, often called Wendland functions [23, 63], to produce a sparse, banded linear systems, which can be solved efficiently. We opt for the second approach: local RBF interpolation using polyharmonic splines and appended polynomial basis functions, which is the focus of this subsection.
@@
Partitions are demonstrated in two examples depicted in Figure 3.1. The Delaunay triangulation is popular for partitioning planar domains and widely integrated into many standard mathematical software libraries [26]. Interpolation nodes are used as the vertices of triangular elements that form the partition (Figure 3.1A). The defining feature of the Delaunay triangulation is that a circle circumscribing each triangular element contains no other interpolation nodes except its vertices. Alternatively, Voronoi partitions (Figure 3.1B) assign each element a single interpolation node and define it as the set of points in $D$ closer to that node than to any other. The Voronoi diagram is the geometric dual of the Delaunay triangulation for a given set of points [^2] This is almost always true. Non-uniqueness can arise when a Voronoi vertex is shared by more than three cells (e.g., in a Cartesian grid), but this is rare for randomly chosen points and can be resolved by small perturbations..
Figure 3.1: Stencils. Each panel shows a subset of the interpolation nodes $\Xi_{n}$ (black dots) for a domain $D\subset\mathbb{R}^{2}$, the boundaries (solid lines) enclosing partition elements $E_{j}$, and an example element (shaded gray) and its stencil points (green circles). A. Delaunay triangulations partition elements using interpolation nodes as vertices. An element’s circumscribing circle does not contain any additional interpolation nodes. B. Voronoi partitions assign one interpolation node per element. A circle centered at a Voronoi vertex can always be drawn to contain only the centers of neighboring elements.
@@
This notion arises naturally both the Delaunay triangulation and Voronoi partition. In the Delaunay case, the radius of any element’s circumscribed circle (See Figure 3.1A) provides a lower bound on $h$, and if the triangulation covers $D$, then $h$ is the maximum of these radii across all triangular elements. In the Voronoi case, the covering radius is the maximum distance from a node to a vertex of its associated patch. Figure 3.1B shows a circle centered on a Voronoi vertex with radius equal to the distance to the adjacent nodes. In both cases, the covering radius is the maximum over a finite set of such radii, making it computationally straightforward.
@@
$s_{E}(\bm{x})=\sum_{\bm{y}\in S(E)}c_{E,\bm{y}}\varphi_{\bm{y}}(\bm{x})+\sum_{|\alpha|\leq\text{deg}}d_{E,\alpha}\pi_{\alpha}(\bm{x}).$
The RBFs $\psi_{\bm{y}}$ are each translates of the same basic function $\psi_{\bm{y}}(\bm{x})=\Phi(\|\bm{x}-\bm{y}\|)$ [23], but not all radially symmetric functions are suitable basic functions. We will restrict ourselves to polyharmonic splines
$\Phi(r)=\begin{cases}r^{\ell},&\text{for $\ell$ odd},\\ r^{\ell}\log(r),&\text{for $\ell$ even.}\end{cases}$ (13)
@@
$f(\bm{x})$ $=s_{E}(\bm{x}),$ for $\bm{x}\in S(E)$, (interpolation)
00 0 $=\sum_{\bm{y}\in S(E)}d_{\alpha}\bm{y}^{\alpha},$ for $|\bm{\alpha}|\leq\text{deg}$. (moment)
Note that though we refer to these as “interpolation conditions”, a point $\bm{x}\in S(E)$ but $\bm{x}\not\in E$ contributes to the interpolant $s_{E}$, but not necessarily the global piecewise interpolant $s$. The full approximation interpolates $f$ at each $\bm{x}\in\Xi$ since every node lies in some element $E$, but each local interpolant $s_{E}$ satisfies additional conditions not directly reflected in the global approximation.
@@
$\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\\ \end{bmatrix}\begin{bmatrix}\bm{c}\\ \bm{d}\end{bmatrix}=\begin{bmatrix}\bm{f}_{E}\\ \bm{0}\end{bmatrix}$ (14)
@@
We must also choose $\ell$, the degree of the polyharmonic spline (see eq. 13). While increasing $\ell$ can reduce approximation error [24], the improvement diminishes quickly. More importantly, polyharmonic splines are conditionally positive definite, and the interpolation matrix eq. 14 is guaranteed to be non-singular if $\text{deg}\leq\lfloor\ell/2\rfloor+1$ [63, 23].

### v2

@@ Abstract
We present a numerical framework for solving neural field equations on surfaces using Radial Basis Function (RBF) interpolation and quadrature. Neural field models describe the evolution of macroscopic brain activity, but modeling studies often overlook the complex geometry of curved cortical domains. Traditional numerical methods, such as finite element or spectral methods, can be computationally expensive and challenging to implement on irregular domains. In contrast, RBF-based methods provide a flexible alternative by offering interpolation and quadrature schemes that efficiently handle arbitrary geometries with high order accuracy. We first develop an RBF-based interpolatory projection framework for neural field models on general smooth surfaces. Quadrature for both flat and curved domains are derived in detail, ensuring high-order accuracy and stability as they depend on RBF hyperparameters (basis functions, augmenting polynomials, and stencil size). Through numerical experiments, we demonstrate the convergence of our method, highlighting its advantages over traditional approaches in terms of flexibility and accuracy. We conclude with an exposition of numerical simulations of spatiotemporal activity on complex surfaces, illustrating the method’s ability to capture complex wave propagation patterns.
@@ 1 Introduction
Neural field equations are nonlinear integro-differential equations that model large-scale neuronal activity, offering tractable formulations for analysis and simulation across diverse domains [13]. Such activity is strongly determined by the architecture of synaptic connections between neurons [20]. Associated models typically grapple with the high volume of neurons and synapses, especially in primate cortex, with approximations such as coarsening, homogenization, and timescale separations [11]. Neural fields treat cortex as a continuum excitable medium and model connectivity via a spatial weight kernel in a nonlinear integral operator, allowing the application of theory from nonlinear waves and other methods from partial differential equations (PDE) [13]. The kernel defines a spatially dependent synaptic coupling strength based on the locations of pre- and post-synaptic neurons. Analyses and simulations typically center on this nonlocal term, which drives local dynamics throughout the domain.
@@
Standard derivations of neural fields are grounded in the known spatial and functional structure of synaptic connectivity throughout cerebral cortex, justifying the typical spatial coarse graining required to obtain continuum models [68, 59]. For analytical and numerical convenience, most modeling studies consider canonical spatial domains and connectivity functions (e.g., distance-dependent weight kernels on planar domains) [21]. Amari (1977) [2] introduced a simplified Heaviside transfer function to enable the use of interface methods and explicit construction of wave solutions. Weight kernels with rational Fourier transforms reduce nonlocal neural fields to local PDEs, facilitating efficient simulation and analysis [40]. However, these simplifications often ignore geometrical complexity. Here we advance a method that is highly flexible to the formulation of neural fields on arbitrary domains with curvature and nonstandard weight kernels.
Although cortex is three-dimensional, many imaging modalities (e.g., fMRI, VSD) access activity on its two-dimensional surface [41, 34]. Prior models considering curved domains appeal to idealizations like spheres [15, 65] and tori [38], allowing for spectral solution methods leveraging known orthonormal eigenbases. General curved surfaces that cannot be expressed as Cartesian products lack this structure. Recent work has explored neural fields on arbitrary smooth surfaces [45], but without convergence guarantees and with only first-order accuracy. A recent framework for projection-based neural field solvers [5] offers an operator-theoretic approach to convergence analysis, decomposing numerical error into contributions from spatial projection, time integration, and quadrature. This unified perspective applies to both Galerkin and collocation methods, clarifying how errors scale with resolution and guiding the design of efficient discretizations for simulations on complex geometries.
In this work, we develop a high-order numerical method for simulating neural field equations on complex geometries and advance the associated convergence theory within a projection-based framework. Our approach combines radial basis function quadrature (RBF-QF) [52, 54, 53] with RBF interpolation to achieve high-order accuracy using relatively few degrees of freedom. This is especially important for simulations on curved surfaces, where evaluating nonlocal integral operators at $\mathcal{O}(n^{2})$ cost can be prohibitive. [^1] On certain domains, FFT-based acceleration reduces complexity to $\mathcal{O}(n\log n)$, but such approaches are not applicable to general manifolds. We formulate projection and collocation schemes for neural fields and present convergence results for projection methods (Section 2); we describe the RBF-QF algorithm and its extension to manifolds (Section 3); and we verify performance through numerical tests and simulations on nontrivial geometries (Sections 4 and 5).
@@ 2 Collocation schemes for the neural field model
Step 1: collocation. Select $n$ points $\Xi=\{\bm{x}_{i}:i\in\mathbb{N}_{n}:=\{1,...,n\}\}\subset D$ in the domain, where $n$ may depend on a discretization parameter $h$ so $n(h)\to\infty$ as $h\to 0$. We impose that Eq. 1 holds at each node, that is, we collocate the equation on $\Xi$. This leads to a functional equation relating $\{\partial_{t}u(t,\bm{x}_{i})\}_{i}$ and $\{u(t,\bm{x}_{i})\}_{i}$ to the function $u(t,{\mkern 2.0mu\cdot\mkern 2.0mu})$, appearing under the integral of Eq. 1.
@@
Henceforth we assume the following hypotheses, provide a natural functional setup for the neural field problem posed on $\mathbb{X}=C(D)$ [51, 5] {hypothesis}[General hypotheses]
@@
To set the stage for our collocation scheme, we rewrite the neural field Eq. 1 in operator form
(2) $u^{\prime}(t)=-u(t)+WF(u(t))+g(t)=:N(t,u(t)),\qquad t\in J\equiv[0,T],$
@@
(3) $W\colon\mathbb{X}\to\mathbb{X},\quad v\mapsto\int_{D}w({\mkern 2.0mu\cdot\mkern 2.0mu},\bm{y})v(\bm{y})\,d\mu(\bm{y}),\qquad F\colon\mathbb{X}\to\mathbb{X},\quad v\mapsto f(v).$
Note, Eq. 1 evolves the $\mathbb{R}$-valued function $u\colon D\times J\to\mathbb{R}$, whereas Eq. 2 evolves the $\mathbb{X}$-valued function $U\colon J\to\mathbb{X}$, $t\mapsto u({\mkern 2.0mu\cdot\mkern 2.0mu},t)$. With a small abuse of notation, we adopt the same symbol for both functions ($u$), as for the forcing function $g$.
Equation 2 describes a Cauchy problem on the Banach space $\mathbb{X}$, which is well-posed under Section 2. That is, there exists a unique $u\in C^{1}(J,\mathbb{X})$ satisfying Eq. 2, as shown in [5, Lemma 2.7] (See [51] for analogous results on unbounded cortices).
2.1 Step 1: collocation via RBF interpolatory projection
In this setup, one seeks a solution $u$ to the neural field equations as a mapping on $[0,T]$ to $\mathbb{X}=C(D)$, the space of continuous functions on $D$. An interpolatory projector $P_{n}$ linearly projects any function $v\in\mathbb{X}$ to a function $v_{n}=P_{n}v$ in an $n$-dimensional subspace $\mathbb{X}_{n}=\operatorname{Span}\{L_{1},\ldots,L_{n}\}\subset\mathbb{X}$, with the property
(4) $P_{n}\colon\mathbb{X}\to\mathbb{X}_{n},\qquad P_{n}v=v,\qquad\text{on $\Xi$.}$
The approximant then takes the form
$v_{n}(\bm{x}):=(P_{n}v)(\bm{x})=\sum_{j=1}^{n}v(\bm{x}_{j})L_{j}(\bm{x}),\qquad x\in D.$
@@ 2.1 Step 1: collocation via RBF interpolatory projection
In view of [5, Proposition 4.1], this is equivalent to imposing that Eq. 1 holds at all $\bm{x}_{i}\in\Xi$ (hence we have collocated the equation, and completed Step 1),
@@
In [5, Theorem 3.1] it is shown that, under Section 2, also the projected problem is well-posed, admitting a unique solution $u_{n}\in C^{1}(J,\mathbb{X}_{n})$.
The projected provlems Eq. 5 evolves the function $u_{n}(t)$ in $\mathbb{X}_{n}$. This form is useful in the analysis of the scheme, but for implementations we pursue an ODE approximating the $n$ coefficients $\alpha_{i}(t):=u_{n}(t)(\bm{x}_{i})$. Since $u_{n}(t)\in\mathbb{X}_{n}$, we set $u_{n}(t)=\sum_{j\in\mathbb{N}_{n}}\alpha_{j}(t)L_{j}$, exploit the Lagrange property of the basis $\{L_{j}\}$, and obtain an ODE in $\mathbb{R}^{n}$
(6) $\alpha^{\prime}(t)=-\alpha(t)+K(\alpha(t))+\gamma(t)\qquad t\in J,$
@@
$\gamma_{i}(t)=g(\bm{x}_{i},t),$ $K_{i}(\alpha)=\int_{D}w(\bm{x}_{i},\bm{y})f\Bigl(\sum_{j\in\mathbb{N}_{n}}\alpha_{j}L_{j}(\bm{y})\Bigr)\,d\mu(\bm{y}).$
@@ 2.2 Step 2: quadrature and Discrete Projection Schemes
The ODE Eq. 6 on $\mathbb{R}^{n}$ is not yet implementable on a computer, because generally the integrals in $K$ can not be evaluated in closed form, and must be approximated by numerical quadrature. Such an approach generates a computationally implementable algorithm using Discrete Projection Schemes.
@@
(7) $W_{n}\colon\mathbb{X}\to\mathbb{X}_{n},\qquad(W_{n}v)(\bm{x})=Q_{n}(w(\bm{x},{\mkern 2.0mu\cdot\mkern 2.0mu})v),$
@@
(8) $\tilde{u}^{\prime}_{n}(t)=-\tilde{u}_{n}(t)+P_{n}W_{n}F(\tilde{u}_{n}(t))+P_{n}g(t),\qquad t\in J,$
@@
Following similar steps to Section 2.1, we obtain
(9) $\tilde{\alpha}^{\prime}(t)=-\tilde{\alpha}(t)+\tilde{K}(\tilde{\alpha}(t))+\gamma(t)\qquad t\in J,$
@@ 2.3 Convergence of the collocation scheme
Under Section 2, if $P_{n}v\to v$ for all $v\in\mathbb{X}$, then $u_{n}\to u$ in $C(J,\mathbb{X})$. Further, there exist positive constants $m$ and $M$, independent of $n$, such that
@@ 3.1 Local Interpolation
The first use of RBF interpolation was by Hardy in 1971 [33] to construct a topography from scattered elevation measurements. The interpolant was a linear combination of radially-symmetric basis functions called Multiquadrics. Critically, the centers of these basis functions coincided with the locations of the measurements, and as a result, the interpolants were guaranteed to exist and be unique. The Mairhuber-Curtis theorem shows that such guarantees break down when basis functions are chosen independently of the interpolation nodes [25, 44, 24] as occurs in multivariate polynomial interpolation. Later developments showed that other radially symmetric basis functions could be used similarly, with resulting interpolants exhibiting spectral convergence as the number of nodes increased  [17, 70]. For a comprehensive introduction to RBF interpolation, see [25].
The price of spectral convergence, however, is that finding interpolant coefficients requires solving an $n\times n$ linear system, which becomes prohibitively expensive and ill-conditioned as $n$ (the number of interpolation nodes) increases. There are two approaches to remedy this limitation. First, one can choose compactly supported basis functions, often called Wendland functions [25, 67], to produce a sparse, banded linear systems, which can be solved efficiently. We opt for the second approach: local RBF interpolation using polyharmonic splines and appended polynomial basis functions, which is the focus of this subsection. For each stencil, unisolvency of the appended polynomial basis is ensured by selecting distinct, well-distributed nodes, so that the associated Vandermonde matrix is full rank. On curved surfaces, we work in local coordinate parameterizations of each stencil, so the same unisolvency conditions as in the planar case apply [10].
RBF interpolation is often touted as a “mesh-free” method and this is certainly true of global RBF interpolation, which yields a smooth interpolant over the entire domain. Local RBF interpolation is arguably mesh-free as well, however it produces a piecewise interpolant that is smooth over each element in a partition of the domain but potentially discontinuous along the boundaries of these elements. Generally, this partition is taken to be the Voronoi partition and is not explicitly calculated. It is thus mesh-free in the sense that the mesh is implicit. We emphasize this because the RBF-QF algorithm for surfaces will explicitly require a triangulation in lieu of a Voronoi partition, and is thus not a mesh-free algorithm. We now provide a self-contained description of the interpolation algorithm.
@@
Partitions are demonstrated in two examples depicted in Figure 1. The Delaunay triangulation is popular for partitioning planar domains and widely integrated into many standard mathematical software libraries [28]. Interpolation nodes are used as the vertices of triangular elements that form the partition (Figure 1A). The defining feature of the Delaunay triangulation is that a circle circumscribing each triangular element contains no other interpolation nodes except its vertices. Alternatively, Voronoi partitions (Figure 1B) assign each element a single interpolation node and define it as the set of points in $D$ closer to that node than to any other. The Voronoi diagram is the graph dual of the Delaunay triangulation for a given set of points [^2] This is almost always true. Non-uniqueness can arise when a Voronoi vertex is shared by more than three cells (e.g., in a Cartesian grid), but this is rare for randomly chosen points and can be resolved by small perturbations..
Figure 1: Stencils. Each panel shows a subset of the interpolation nodes $\Xi_{n}$ (black dots) for a domain $D\subset\mathbb{R}^{2}$, the boundaries (solid lines) enclosing partition elements $E_{j}$, and an example element (shaded gray) and its stencil points (green circles). A. Delaunay triangulations partition elements using interpolation nodes as vertices. An element’s circumscribing circle does not contain any additional interpolation nodes. B. Voronoi partitions assign one interpolation node per element. A circle centered at a Voronoi vertex can always be drawn to contain only the centers of neighboring elements.
@@
This notion arises naturally in both the Delaunay triangulation and Voronoi partition. In the Delaunay case, the radius of any element’s circumscribed circle (See Figure 1A) provides a lower bound on $h$, and if the triangulation covers $D$, then $h$ is the maximum of these radii across all triangular elements. In the Voronoi case, the covering radius is the maximum distance from a node to a vertex of its associated patch. Figure 1B shows a circle centered on a Voronoi vertex with radius equal to the distance to the adjacent nodes. In both cases, the covering radius is the maximum over a finite set of such radii, making it computationally straightforward.
@@
$s_{E}(\bm{x})=\sum_{\bm{y}\in S(E)}c_{E,\bm{y}}\varphi_{\bm{y}}(\bm{x})+\sum_{|\bm{\alpha}|\leq\text{deg}}d_{E,\bm{\alpha}}\pi_{\bm{\alpha}}(\bm{x}).$
The RBFs $\psi_{\bm{y}}$ are each translates of the same basic function $\psi_{\bm{y}}(\bm{x})=\Phi(\|\bm{x}-\bm{y}\|)$ [25], but not all radially symmetric functions are suitable basic functions. We will restrict ourselves to polyharmonic splines with fixed order parameter $\ell$
(12) $\Phi(r)=\begin{cases}r^{\ell},&\text{for $\ell$ odd},\\ r^{\ell}\log(r),&\text{for $\ell$ even.}\end{cases}$
@@
(interpolation) $f(\bm{x})$ $=s_{E}(\bm{x}),$ for $\bm{x}\in S(E)$,
(moment) $0$ $=\sum_{\bm{y}\in S(E)}d_{\alpha}\bm{y}^{\alpha},$ for $|\bm{\alpha}|\leq\text{deg}$.
Note that though we refer to these as “interpolation conditions”, a point $\bm{x}\in S(E)$ but $\bm{x}\not\in E$ is interpolated by the local interpolant $s_{E}$, but not by the global piecewise interpolant $s$. The full approximation interpolates $f$ at each $\bm{x}\in\Xi$ since every node lies in some element $E$, but each local interpolant $s_{E}$ satisfies additional conditions not directly reflected in the global approximation.
@@
(13) $\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\\ \end{bmatrix}\begin{bmatrix}\bm{c}\\ \bm{d}\end{bmatrix}=\begin{bmatrix}\bm{f}_{E}\\ \bm{0}\end{bmatrix}$
@@
We must also choose $\ell$, the degree of the polyharmonic spline (see Eq. 12). Although increasing $\ell$ can reduce error, the improvement diminishes quickly, and non-singularity of the interpolation matrix requires $\text{deg}\geq\lfloor\ell/2\rfloor+1$ [67, 25]. In practice, we therefore adopt parameters that balance accuracy, stability, and cost: stencil sizes $k=21$ or $32$, slightly larger than the minimum needed for unisolvency, improve stability near boundaries without substantially increasing cost [26, 9]; and we fix $\ell=3$, since larger values yield little accuracy gain but worsen conditioning [67]. With appended polynomials up to degree $\deg=4$, these choices consistently performed well, and convergence tests (Sections 3.23.3) confirm robust accuracy across both flat and curved geometries.